# GRIP: A General Robotic Incremental Potential Contact Simulation Dataset for Unified Deformable-Rigid Coupled Grasping

### v1

@@ I INTRODUCTION
Grasping is pivotal to robotics, serving as the cornerstone of nearly all manipulation tasks.
However, acquiring reliable grasping data for robotics training faces challenges such as object variability, occlusions, and generalization to unseen objects. The emergence of large-scale grasp datasets, such as GraspNet-1Billion [1] and DexGraspNet [2], has significantly advanced research in this area, providing rich data to improve grasp prediction, facilitating applications across industrial automation, household robotics, and healthcare.
@@
The recently proposed Incremental Potential Contact (IPC) method [6] serves as a remedy, effectively addressing large-deformation and frictional contact problems for soft bodies within a variational framework, while guaranteeing intersection- and inversion-free solutions. Recent extensions and developments, such as [7, 8, 9, 10, 11], have integrated IPC and affine body dynamics (ABD) into a unified framework. However, these approaches either lack user-friendly interfaces for robotics, or are not optimized for large-scale, multi-environment simulations, resulting in inefficiencies in robotic applications.
@@ II-A Dataset for Hand Grasping
Existing grasp datasets for robotic grippers focus mainly on interactions between rigid grippers and rigid objects [14, 15, 16, 17]. Such grippers consist of rigid links connected through joints.
For example, [1, 18] provide grasp poses for rigid parallel grippers. Dexterous rigid grippers, which offer greater agility and versatility due to their increased degrees of freedom (DoFs), enable a wider range of human-like grasping motions. [2] synthesizes dexterous poses for ShadowHand, whereas [19] generates grasp poses for five different dexterous hands. More recently, [20] introduced a large-scale synthetic dataset covering 11 common types of rigid grippers. In addition to single grippers, [21] introduces BimanGrasp-Dataset, a large-scale synthetic dataset focusing on bimanual grasping with heavy or large objects.
@@ II-B Grasp Generation
Recently, data-driven methods [26, 27, 28, 29, 30] have gained significant attention. These approaches predict grasp poses based on object geometry and efficiently generate diverse grasps in parallel, making them well-suited for real-time applications. Moreover, many studies [26, 28, 29, 30] support point cloud representations of grasped objects, unlike most optimization-based methods, further improving their practicality for real-world deployment.
@@ II-C Robot Simulation
Mainstream robot simulation algorithms for deformable bodies primarily adopt either the Material Point Method (MPM) [31, 32, 33] or the Finite Element Method (FEM) [34, 35, 10, 36]. MPM models an object as a set of particles without explicit topology, effective for handling large topology changes. However, it relies on a background Eulerian grid, which introduces numerical diffusion artifacts and incurs high computational costs as grid resolution increases [37, 38]. FEM represents objects as structured meshes, such as triangular elements for surface meshes and tetrahedral elements for volume meshes, and yields great accuracy when objects undergo small deformations. Nevertheless, for most FEM-based methods, large deformations will lead to element inversion, resulting in simulation failures [39, 40]. Further, unlike MPM, FEM requires explicit collision handling, which is commonly formulated as a Linear Complementarity Problem (LCP) [41, 42] or a convex optimization problem [43, 44]. LCP-based methods are hindered by their NP-hardness and ill-conditioning issues, while convex-optimization-based approaches face challenges in extreme cases, such as fast-moving non-convex objects and extremely thin objects.
Recently, Li et al. proposed Incremental Potential Contact (IPC) [6], a FEM-based method that accurately and robustly handles extreme contact scenarios. IPC formulates frictional contact within FEM as an optimization problem, incorporating barrier augmentation, continuous collision detection (CCD), and a Projected Newton optimization scheme with a step-size filter to enforce intersection- and inversion-free constraints [45]. Despite its high accuracy in robotic simulations [46, 47, 48], the computational inefficiency of IPC limits its applicability to large-scale simulations, hindering its adoption in large data generation and parallel reinforcement learning (RL) training.
Our work extends IPC to large-scale parallel environment simulation, improving its efficiency and scalability through parallel computing. We construct a large-scale grasp pose dataset using optimization-based pose synthesis combined with a multi-environment parallel IPC simulator.
@@ III-B Grippers and Objects
Our dataset covers both UMI grippers [3] and LEAP Hands [12]. For simulation, we tetrahedralize the soft finger meshes of UMI grippers using TetWild [49], setting a tight envelope tolerance of $\epsilon_{e}=5\times 10^{-4}$ to preserve fine surface geometry details. The Young’s modulus ($E_{\text{UMI}}=9.4\times 10^{6}$ Pa) and dry friction coefficient ($\mu_{\text{UMI}}=3.5$) of the UMI gripper fingers are set to match their real-world counterparts [3].
For LEAP Hands, leveraging IPC’s robustness for handling complex geometries,
we use their high-resolution visual meshes directly as collision meshes for precise contact resolution in simulation. To generate the contact point candidates for DexGraspNet synthesis, we heuristically sample points evenly across the contact regions of the hand’s palms and fingers.
@@
We select 800 objects from the DexGraspNet dataset [2] for single-hand grasping. For bimanual grasping, we retain 400 large objects from the single-hand set and add additional 400 objects from the PartNet-Mobility dataset [50], totaling 800 large-scale objects. To prepare for simulation, we preprocess the surface meshes to be watertight and tetrahedralize them for soft body simulations using TetWild. We also apply domain randomization, sampling Young’s modulus, friction coefficient, and object scaling to enhance data diversity.
@@ III-D Grasp Validation
At the start of each trajectory, gravity is disabled to prevent the object from falling before being grasped. To simulate contact-aware grasping, a finger’s motion is halted once its contact force exceeds 50 N. After the system reaches a steady state, gravity is applied in six axis-aligned directions, each with a magnitude of 9.8 m/$\text{s}^{2}$ for 0.1 seconds. A grasp pose is considered stable if the object remains in contact with the gripper at the final time step of the trajectory and if its center of mass moves less than a threshold distance proportional to $\epsilon_{v}\Delta t.$
@@ IV-B Real2Sim Validation
We select four objects of varying sizes and shapes from the single-hand manipuland set and fabricate them using a 3D printer with TPE-83A material. These objects are then grasped by a UMI gripper, and the resulting trajectories are recorded with a calibrated RealSense D435 camera. ArUco markers are attached to the manipulands and the grippers for 6D pose detection [51]. The recorded gripper trajectories are subsequently replayed in our simulator, with the initial poses of the manipulands aligned to those in real-world setup. As shown in Fig. 4, the bending behavior of the UMI gripper fingers and the 6D poses of the manipulands in the simulation closely match the real-world results, demonstrating high physical accuracy and fidelity.
@@ IV-C Dataset Quality Evaluation
To quantitatively analyze the grasp pose quality, we employ two metrics: (1) penetration distance $D_{1}$, which measures the penetration depth between the object and the gripper, if any, and (2) absolute distance $D_{2}$, defined as the unsigned distance between the object and the gripper to assess the grasp tightness. Fig. 6 illustrates the computation of these two metrics.
Both metrics are computed using a Signed Distance Field (SDF) $d^{\text{o}}:~{}\mathbb{R}^{3}\rightarrow\mathbb{R}$ constructed for the manipuland, with point-SDF distances computed by sampling 50,000 points (forming a set $S^{\text{h}}$) on the LEAP Hand’s surface:
@@
Using these two metrics, we evaluate the grasp quality produced by two simulators on a bottle subset (111 bottles) and the full 800-objects dataset. As shown in Table I, our method achieves zero penetration cases, whereas Isaac Gym exhibits significant penetration artifacts (see Fig. 5), with an average penetration depth of 5 mm on the bottle subset and 8.7 mm across the full dataset. Additionally, the absolute distance in our method closely aligns with the IPC distance threshold parameter ($\hat{d}=10^{-3}$ m) used for surface compliance, indicating that the resulting grasps are tightly secured. In contrast, poses generated with Issac Gym exhibit significantly larger absolute distances, suggesting that the gripper may fail to hold the manipuland firmly. These results demonstrate that our IPC-based simulator not only eliminates penetration artifacts but also achieves stable grasps, making it a more reliable choice for generating high-quality grasping datasets.
@@ IV-D Stress Force Prediction
To further assess the accuracy and effectiveness of our stress prediction network, we apply the real-to-sim pipeline described in Section III-D and conduct four grasping experiments using 3D-printed mug and bowl objects. Fig. 8 shows the normalized stress distribution predicted by our network alongside the simulation results, demonstrating strong agreement. For the grasp pose shown in the first row, the bowl undergoes significant deformation; rather, the grasp pose in the second row yields a moderate deformation. This aligns well with our network’s predictions, where the mean stress in the first case is $2.17$ times that in the second. Similarly, for the two mug grasps, the predicted mean stress of the former is 1.6 times that of the latter, in accordance with visual deformation. Both the qualitative and the quantitative evaluations confirm the high accuracy of our stress prediction network.
@@ V CONCLUSION
We introduce an IPC-based simulator optimized for large-scale, highly efficient, parallel multi-environment simulations, addressing key challenges in simulating soft gripper and deformable object grasping. Additionally, we develop a fully automated pipeline for grasp generation, simulation, and evaluation, supporting a wide range of grippers and manipuland types. Enabled by this pipeline, we construct the GRIP dataset, a large-scale, diverse grasp dataset featuring both soft and rigid grippers, to advance research in neural grasp generation and soft object manipulation.
Experimental results demonstrate the effectiveness of our pipeline in synthesizing stable grasp poses and accurately predicting manipuland stress distributions. Through a real-to-simulation comparison, we further validate the strong agreement between our simulated results and real-world physics. However, our current grasp synthesis does not incorporate simulation feedback, as grasp generation is performed independently of the resulting dynamics. In future work, we aim to integrate differentiability into our pipeline to establish a closed-loop system, enabling dynamic feedback to refine grasp synthesis.

### v2

@@ I INTRODUCTION
Grasping is pivotal to robotics, serving as the cornerstone of nearly all manipulation tasks. However, acquiring reliable grasping data for robotics training faces challenges such as object variability, occlusions, and generalization to unseen objects. The emergence of large-scale grasp datasets, such as GraspNet-1Billion [1] and DexGraspNet [2], has significantly advanced research in this area, providing rich data to improve grasp prediction, facilitating applications across industrial automation, household robotics, and healthcare.
@@
The recently proposed Incremental Potential Contact (IPC) method [6] serves as a remedy, effectively addressing large-deformation and frictional contact problems for soft bodies within a variational framework, while guaranteeing intersection- and inversion-free solutions. Recent extensions and developments, such as [7, 8], have integrated IPC and Affine Body Dynamics (ABD) [9] into a unified framework with multibody and articulated dynamics [10]. However, these approaches either lack user-friendly interfaces for robotic applications, or are not optimized for large-scale, multi-environment simulations, resulting in inefficiencies in robotic applications.
@@ II-A Dataset for Hand Grasping
Existing grasp datasets for robotic grippers primarily focus on rigid gripper–rigid object interactions. Many center on parallel grippers [13, 14], object-centric relationships [15], or multimodal sensing [16]. These grippers are typically composed of rigid links connected via joints. For example, [1, 17] provide grasp poses for rigid parallel grippers. Dexterous rigid grippers, which offer greater agility and versatility due to their increased degrees of freedom (DoFs), enable a wider range of human-like grasping motions. [2, 18] synthesize dexterous poses for ShadowHand and Allegro, whereas [19] generates grasp poses for five different dexterous hands. [20] introduced a large-scale synthetic dataset covering 11 common types of rigid grippers. In addition to single grippers, [21] introduces a large-scale synthetic dataset focusing on bimanual grasping with heavy or large objects.
@@ II-B Grasp Generation
Recently, data-driven methods have advanced grasp generation through diffusion-based dexterous grasping [26], reinforcement learning [27], and grasp proposal networks for both 6-DOF [28, 29] and dexterous grasping [30]. These approaches predict grasp poses based on object geometry and efficiently generate diverse grasps in parallel, making them well-suited for real-time applications. Moreover, many studies [26, 28, 29, 30] support point cloud representations of grasped objects, further improving their practicality for real-world deployment.
@@ II-C Robot Simulation
Mainstream robotic deformable body simulation algorithms often adopt the Material Point Method (MPM) for tasks such as tactile sensing [31], soft robot control [32], and soft object manipulation [33], or use the Finite Element Method (FEM) [34, 35] with various dimensional topologies [8, 36]. MPM models an object as a set of particles without explicit topology, effective for handling large topology changes. However, it relies on a background Eulerian grid, which introduces numerical diffusion artifacts and incurs high computational costs as grid resolution increases [37, 38]. FEM represents objects as structured meshes, such as triangular elements for surface meshes and tetrahedral elements for volume meshes, and yields great accuracy when objects undergo small deformations. Nevertheless, for most FEM-based methods, large deformations will lead to element inversion, resulting in simulation failures [39, 40]. Further, unlike MPM, FEM requires explicit collision handling, which is commonly formulated as a Linear Complementarity Problem (LCP) [41, 42] or a convex optimization problem [43, 44]. LCP-based methods are hindered by their NP-hardness and ill-conditioning issues, while convex-optimization-based approaches face challenges in extreme cases, such as fast-moving non-convex objects and extremely thin objects.
Recently, Li et al. proposed Incremental Potential Contact (IPC) [6], a FEM-based method that accurately and robustly handles extreme contact scenarios. IPC formulates frictional contact within FEM as an optimization problem, incorporating barrier augmentation, continuous collision detection (CCD), and a Projected Newton optimization scheme with a step-size filter to enforce intersection- and inversion-free constraints [45]. Despite its high accuracy in robotic simulations, including dataset generation [46], tactile modeling [47], and general-purpose simulation [48], the computational inefficiency of IPC limits its applicability to large-scale simulations, hindering its adoption in large data generation and parallel reinforcement learning training.
Our work extends IPC to a large-scale parallel environment simulation, improving its efficiency and scalability through parallel computing. We construct a large-scale grasp pose dataset using optimization-based pose synthesis combined with a multi-environment parallel IPC simulator.
@@ III-B Grippers and Objects
Our dataset covers both UMI grippers [3] and LEAP Hands [11]. For simulation, we tetrahedralize the soft finger meshes of UMI grippers using TetWild [49], setting a tight envelope tolerance of $\epsilon_{e}=5\times 10^{-4}$ to preserve fine surface geometry details. The Young’s modulus ($E_{\text{UMI}}=9.4\times 10^{6}$ Pa) and dry friction coefficient ($\mu_{\text{UMI}}=3.5$) of the UMI gripper fingers are set to match their real-world counterparts [3]. For LEAP Hands, leveraging IPC’s robustness for handling complex geometries, we use their high-resolution visual meshes directly as collision meshes for precise contact resolution in simulation. To generate the contact point candidates for DexGraspNet synthesis, we heuristically sample points evenly across the contact regions of the hand’s palms and fingers.
@@
We select 800 objects from the DexGraspNet dataset [2] for single-hand grasping via stratified sampling based on label categories and their object counts. The subset matches the original distribution, with at least one object included from underrepresented labels to ensure diversity. For bimanual grasping, we retain 400 large objects from the single-hand set and add additional 400 objects from the PartNet-Mobility dataset [50], totaling 800 large-scale objects. To prepare for the simulation, we preprocess the surface meshes to be watertight and tetrahedralize them for soft body simulations using TetWild. We also apply domain randomization, sampling Young’s modulus, friction coefficient, and object scaling to enhance data diversity.
@@ III-D Grasp Validation
At the start of each trajectory, gravity is disabled to prevent the object from falling before being grasped. To simulate contact-aware grasping, a finger’s motion is halted once its contact force exceeds 50 N. After the system reaches a steady state, gravity is applied in six axis-aligned directions, each with a magnitude of 9.8 m/$\text{s}^{2}$ for 0.1 seconds. A grasp pose is considered stable if the object remains in contact with the gripper at the final time step of the trajectory and if its center of mass moves less than a threshold proportional to $\epsilon_{v}\Delta t.$
@@ IV-B Real2Sim Validation
We select four objects of varying sizes and shapes from the single-hand manipuland set and fabricate them using a 3D printer with TPE-83A material. These objects are then grasped by a UMI gripper, and the resulting trajectories are recorded with a calibrated RealSense D435 camera. ArUco markers are attached to the manipulands and the grippers for 6D pose detection [51]. The recorded gripper trajectories are subsequently replayed in our simulator, with the initial poses of the manipulands aligned to those in the real-world setup. As shown in Fig. 4, the bending behavior of the UMI gripper fingers and the 6D poses of the manipulands in the simulation closely match the real-world results, demonstrating high physical accuracy and fidelity.
@@ IV-C Dataset Quality Evaluation
To quantitatively analyze the grasp pose quality, we employ two metrics: (1) penetration distance $D_{1}$, which measures the penetration depth between the object and the gripper, if any, and (2) absolute distance $D_{2}$, defined as the unsigned distance between the object and the gripper to assess the grasp tightness. Fig. 6 illustrates the computation of these two metrics. Both metrics are computed using a Signed Distance Field (SDF) $d^{\text{o}}:~{}\mathbb{R}^{3}\rightarrow\mathbb{R}$ constructed for the manipuland, with point-SDF distances computed by sampling 50,000 points (forming a set $S^{\text{h}}$) on the LEAP Hand’s surface:
@@
Using these two metrics, we evaluate the grasp quality produced by two simulators on a bottle subset (111 bottles) and the full 800-object dataset. As shown in Table I, our method achieves zero penetration cases, whereas Isaac Gym exhibits significant penetration artifacts (see Fig. 5), with an average penetration depth of 5 mm on the bottle subset and 8.7 mm across the full dataset. Additionally, the absolute distance in our method closely aligns with the IPC distance threshold parameter ($\hat{d}=10^{-3}$ m) used for surface compliance, indicating that the resulting grasps are tightly secured. In contrast, poses generated with Issac Gym exhibit significantly larger absolute distances, suggesting that the gripper may fail to hold the manipuland firmly. These results demonstrate that our IPC-based simulator not only eliminates penetration artifacts but also achieves stable grasps, making it a more reliable choice for a high-quality grasping dataset.
@@ IV-D Stress Force Prediction
To further assess the accuracy and effectiveness of our stress prediction network, we apply the real-to-sim pipeline described in Section III-D and conduct four grasping experiments using 3D-printed mug and bowl objects. Fig. 8 shows the normalized stress distribution predicted by our network alongside the simulation results, demonstrating strong agreement. For the grasp pose shown in the first row, the bowl undergoes significant deformation; rather, the grasp pose in the second row yields a moderate deformation. This aligns well with our network’s predictions, where the mean stress in the first case is $2.17$ times that in the second. Similarly, for the two mug grasps, the predicted mean stress of the former is 1.6 times that of the latter, in accordance with visual deformation. Both the qualitative and quantitative evaluations confirm the high accuracy of our stress prediction network.
@@ V CONCLUSION
We introduce an IPC-based simulator optimized for large-scale, highly efficient, parallel multi-environment simulations, addressing key challenges in simulating soft grippers and deformable object grasping. Additionally, we develop a fully automated pipeline for grasp generation, simulation, and evaluation, supporting a wide range of grippers and manipuland types. Enabled by this pipeline, we construct the GRIP dataset, a large-scale, diverse grasp dataset featuring both soft and rigid grippers, to advance research in neural grasp generation and soft object manipulation.
Experimental results demonstrate the effectiveness of our pipeline in synthesizing stable grasp poses and accurately predicting manipuland stress distributions. Our current GRIP dataset features only UMI and LEAP Hands, but our gripper-agnostic pipeline can be easily extended to support diverse soft and rigid grippers. Through a real-to-simulation comparison, we further validate the strong agreement between our simulated results and real-world physics. However, our current grasp synthesis does not incorporate simulation feedback, as grasp generation is performed independently of the resulting dynamics. In future work, we aim to integrate differentiability into our pipeline to establish a closed-loop system, enabling dynamic feedback to refine grasp synthesis.